A manufacturer produces 25-pound lifting weights. The lowest actual weight is 24 pounds, and the highest is 26 pounds. Each weight is equally likely so the distribution of weights is uniform. A sample of 100 weights is taken.

1. What is the distribution for the mean weight of 100 25-pound lifting weights?

2. Find the probability that the mean actual weight for the 100 weights is less than 24.9.

3. Find the probability that the mean actual weight for the 100 weights is greater than 25.2.

4. Find the 90th percentile for the mean weight for the 100 weights.

Please show answer using Excel functions/formula's
Answer #1

a = lower limit = 24

b = upper limit = 26

1.

mean = (a+b)/2 = (24+26)/2 = 25

mean = 25 pound

2.

P(x<24.9) = (24.9 - a)/(b-a)

= (24.9 - 24)/(26-24)

= (0.9)/2 = 0.45

P(x<24.9) = 0.45

3.

P(x>25.2) = (b-25.2) / (b-a)

= (26-25.2)/(26-24)

= 0.8/2

= 0.4

P(x>25.2) = 0.4

4.

P(x<X) = (X-a)/(b-a)

= (X-24)/(26-24)

= (X-24)/2

(X-24)/2 = 0.9

X = 25.8

(please UPVOTE)